OpenStudy (anonymous):

SOMEONE HELP ME WITH THESE PROBLEMS 1. 9.25 x 10^21 pennies is how many moles? 2. 2.68 moles of cobalt atoms is how many total atoms? 3. 1.08 x 10^25 carbon dioxide molecules is how many moles of carbon dioxide? 4. How many molecules does 0.773 moles of Ca(OH)2 contain? 5. 2.13 x 10^24 molecules of water is how many moles of water? 6. How many total atoms are there in 2.21 moles of magnesium bromide? 7. Counting one number per second, how many years would it take you to count from 1 to Avogadro's number?

OpenStudy (aaronq):

you need to use these formulas: \(n=\dfrac{N}{N_A}\) and \(n=\dfrac{m}{M}\) n=moles; N=# of particles; \(N_A\)=avogadros number; M=molar mass; m=mass (in grams)